Hello! I'm having problems with substitution and for the following problem i keep getting X=7 y=4 and z=1. I thought this made sense but my answer sheet says it should be x=2 y=1 and z=1. Could you show me how its done so i can compare it to the way i've been trying?
here's the equation and thank you!!!
-3x-3y+2z=-7
z=1
-2x-3y+z=-6

Answer by stanbon(75887)

-3x-3y+2z=-7
z=1
-2x-3y+z=-6
----------------
Substitute z for z and rewrite the equations:
-3x-3y=-9
-2x-3y=-7
------------
Subtract the 1st equation by the 2nd to get:
x = 2
-------
Substitute into -2x-3y=-7 to solve for y:
-2*2-3y = -7
-3y = -3
y = 1
================
Solution:
x = 2; y = 1, z = 1
==============
Cheers,
Stan H.